# Towards Better Benchmark Datasets for Inductive Knowledge Graph Completion

### v1

@@ 1 Introduction
Knowledge Graph Completion (KGC) attempts to predict unseen facts on an existing knowledge graph (KG). KGC has numerous applications including drug discovery [1], personalized medicine [2], and recommendations [3]. Traditionally, most research on KGC has focused on the transductive setting, where the same entities and relations are seen during training and testing. Most methods [4, 5, 6] generally focus on learning embeddings for all entities and relations to facilitate the prediction of new facts.
In recent years, interest in KGC has shifted towards designing methods that can generalize to new entities or relations not seen during training. This task, known as “inductive KGC”, requires a method to train on a graph $\mathcal{G}_{\text{train}}$ and perform inference on a different graph $\mathcal{G}_{\text{inf}}$, where the inference graph contains either new entities, relations, or both. Because of this, methods for inductive KGC don’t rely on fixed embeddings for entities or relations, instead opting for more flexible techniques that can inductively learn representations based on a given graph [7, 8, 9]. To asses the ability of methods for this task, new datasets have been constructed that require the ability to reason inductively. All inductive datasets [7, 10, 9] are constructed from existing transductive KGC datasets. This is done by sampling two graphs that contain disjoint entities, one each for train and inference. Multiple methods [8, 11, 9] have reported tremendous promise on these newer benchmark datasets.
@@
This finding naturally motivates us to ask – why can PPR perform so well on existing inductive datasets? In Section 3 we discover that the high performance of PPR is due to how the inductive datasets are created from transductive datasets. Specifically, we observe that the current procedure creates graphs where the shortest path distance (SPD) between entities in positive test samples is much lower than the SPD between those in negative samples. This allows for a method like PPR, which gives a higher weight to shorter walks, to distinguish between the positive and negative samples based solely on the distance. To account for this problem, in Section 4 we propose a new strategy for sampling inductive datasets that uses graph partitioning to create the train and inference graphs. This allows us to sample subgraphs that retain the general properties of the original graph. We then demonstrate that this procedure can indeed create inductive datasets that limit the performance ability of PPR. Lastly, we benchmark common inductive KGC methods on our newly constructed datasets. Our key contributions can be summarized as follows:
@@
  • Through empirical study, we find that the strong performance of PPR on inductive datasets is due to the current procedure for constructing those datasets, which allows for a shortcut to distinguish between positive and negative samples.
  • We propose a new strategy for sampling inductive KGC datasets from their transductive counterparts that uses graph partitioning. We show that our proposed method can substantially mitigate the shortcut. We then benchmark popular inductive KGC method on our newly created datasets. Compared to the older datasets, previous SOTA tend to decrease in performance.
The remaining paper is structured as follows. In Section 2, we provide background on inductive KGC methods, datasets, and PPR. In Section 3, we study in detail when and why PPR can perform well on inductive KGC. We then introduce our new strategy for creating inductive datasets in Section 4 and benchmark popular methods on these newly created datasets in Section 5.
@@ 2 Background
Throughout this study we denote a knowledge graph as $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{E}\}$ where $\mathcal{V}$ is a set of entities (i.e., nodes), $\mathcal{R}$ the set of relations (i.e., edge types), and $\mathcal{E}$ the set of edges (i.e., triples) of the form $(s,r,o)$ where $s$ and $o$ are entities and $r$ a relation. Lastly, we note that the task of KGC is formulated as the following where given a query $(s,r,*)$, we attempt to predict the correct entity *.
@@
Inductive KGC Datasets: In inductive KGC, we are given a training graph $\mathcal{G}_{\text{train}}=\{\mathcal{V}_{\text{train}},\mathcal{R}_{\text{train}},\mathcal{E}_{\text{train}}\}$ and an inference graph $\mathcal{G}_{\text{inf}}=\{\mathcal{V}_{\text{inf}},\mathcal{R}_{\text{inf}},\mathcal{E}_{\text{inf}}\}$. A method is trained on $\mathcal{G}_{\text{train}}$ and evaluated on $\mathcal{G}_{\text{inf}}$. Most datasets consider the setting that is only disjoint on the entities such that $\mathcal{R}_{\text{inf}}\in\mathcal{R}_{\text{train}}$ and $\mathcal{E}_{\text{inf}}\cap\mathcal{E}_{\text{train}}=\emptyset$. This setting is referred to as the (E) setting. Another setting, which we denote as (E, R) further allows $\mathcal{R}_{\text{inf}}$ to contain relations not in $\mathcal{R}_{\text{train}}$.
@@
All existing inductive datasets are sampled from existing transductive datasets. Furthermore, the majority of inductive datasets [7, 9] are further created via the same procedure introduced by [7]. We give brief overview of this procedure below. Given a transductive dataset, which we denote as $\mathcal{G}$, $k$ seed entities are randomly chosen from $\mathcal{G}$. The 2-hop neighborhood is then extracted around each individual seed entity. To prevent exponential growth, the number of neighbors sampled at any hop is capped at 50 for each individual seed entity. The resulting edges are then combined to create $\mathcal{G}_{\text{train}}$ and are subsequently removed from the original graph. The inference graph is then sampled in a similar manner from the resulting graph $\mathcal{G}\setminus\mathcal{G}_{\text{train}}$. We note that the one exception to this procedure are the ILPC datasets, introduced by [10], which instead sample $p\%$ of the nodes from $\mathcal{G}$ and use them to create $\mathcal{G}_{\text{train}}$. The rest of the nodes are then used to construct $\mathcal{G}_{\text{inf}}$. In practice, we find that both sampling methods tend to produce similar graphs.
@@
Inductive KGC Methods: Both NeurlLP [13] and DRUM [14] consider combining the path representations of different length between both entities in a triple. However, since they explicitly consider each path, they are often limited to only considering paths of up to length 2 or 3. Conditional MPNNs [15] are a more efficient alternative to encoding higher-order path information. They work by conditioning the message passing mechanism on the known entity, allowing the implicit aggregation of all paths up to a length $L$ (which is equal to the number of GNN layers). The value of $L$ is typically $5/6$. Prominent examples include NBFNet [8] and RED-GNN [11]. More scalable alternatives have been proposed that prune the propagated messages [16, 17, 18]. NodePiece [19] is concerned with parameter efficiency, using an anchor-based approach to learn a more compact set of entity and relation representations. All previous methods assume that a fixed set of relations exist between the train and inference graphs. To account for new relations in inference, InGRAM [9] introduces the concept of a “relation graph”, which inductively encodes the representation of each relation. Note that we omit subgraph methods that have been to shown to prohibitively expensive [7, 20, 21, 22, 23] or those that require the use of external textual information [24].
@@
where $W$ is the random walk matrix and $x_{s}$ is a one-hot vector at node $s$. Intuitively, we can see that weight given to a walk decays with the increase in length due to $(1-\alpha)^{k}$. As such, the PPR score will often be higher for those nodes of shorter distance to $s$. Note that for a KG, we obtain the PPR matrix by first converting the inference graph, $\mathcal{G}_{\text{inf}}$, to an undirected graph. We note that this is common practice in KGC [26] whereby inverse relations are added to the graph. We further assume an edge weight of 1 for all edges.
@@ 3.1 Personalized Pagerank (PPR) is a Strong Baseline for Inductive KGC
We begin by obtaining the PPR matrix of the inference graph for use in evaluation. See Section 2 for more details on how this is done. Given this new graph, for a query $(s,r,*)$, we calculate the PPR score for all possible entities $o\in\mathcal{V}$. Using these scores, we can obtain the rank of the true entity for the query and compute the performance in terms of Hits@10.
In Figure 1, we show the performance when using the PPR versus the SOTA performance among supervised methods. We split the datasets by transductive, (E), and (E, R) inductive. We include those datasets most often used in each task, comprising 25 in total. See Appendix B.1 for more details on the datasets chosen. The full set of results can be found in Appendix A.1. We observe that on both types of inductive datasets, the PPR score does reasonably well, performing on average only 25-29% less than SOTA. Also, on some datasets like the WN or ILPC inductive splits, the performance nearly matches or exceeds the supervised performance. On the other hand, for the transductive datasets, the performance disparity is often much larger. Interestingly, PPR still performs well on some transductive datasets, including the popular WN18RR. This tells us that this phenomenon is not unique to inductive datasets. However, this phenomenon is far less common on transductive datasets.
Curiously, for the inductive datasets, their performance using PPR is much higher than their transductive parents. For example, on the inductive datasets derived from FB15k-237, the performance of PPR increases by an average of 1,186%. Furthermore, for those derived from WN18RR, they see an average increase of 43%. This suggests that there is some change in the underlying inductive graphs that are causing the performance of PPR to increase.
@@ 3.2 When Does PPR Perform Well and Why?
Figure 2: $\Delta\text{SPD}$ vs. PPR Hits@10.
We find that the answer lies in comparing the mean shortest path distance (SPD) between positive and negative samples. For a query $(s,r,*)$, let $o^{+}$ denote the correct answer and $o^{-}\in\mathcal{V}^{-}(s,r)$ as the set of negative answers for the query. We compute the SPD on the inference graph for both $o^{+}$ and $\mathcal{V}^{-}(s,r)$. This is repeated for each test sample $(s,q,*)\in\mathcal{E}_{\text{inf}}$. We then calculate the mean SPD across all positive and negative test samples, which we denote as $\text{SPD}^{+}$ and $\text{SPD}^{-}$, respectively. The different in mean SPD is correspondingly given by $\Delta\text{SPD}=\text{SPD}^{+}-\text{SPD}^{-}$. This tells us, on average, how much shorter the distance between entities in positive samples are relative to those in negatives.
In Figure 2 we show the relationship between $\Delta\text{SPD}$ and the Hits@10 when using PPR. The results are across 25 different datasets, for both inductive and transductive KGC. Calculating the Pearson correlation, we can observe it is 0.87, which indicates a strong relationship between the two metrics. Therefore, there exists a basic pattern to distinguish positive and negative samples in many KGC datasets. Put simply, the SPD between entities in positive test samples tend to be lower than that in negative samples. The larger the discrepancy between the mean distances, the better PPR can perform. This suggests that the PPR scores can exploit this distance disparity in the datasets to achieve good performance, even while completely ignoring the relational information.
But, why can PPR exploit this pattern? As shown in Eq. (1), the PPR score between two nodes is the weighted sum of random walk probabilities between them. Furthermore, walks of shorter length are weighed more heavily than those of longer length. Therefore, nodes with a shorter distance between them will be able to benefit from these highly-weighted walks, while those of larger distance will not. For example, when $\alpha=0.15$, the weight of a walk with $\text{SPD}=2$ is $0.72$ while when $\text{SPD}=5$ it is $0.44$. As such, the PPR score will invariably favor those node pairs with a lower SPD.
@@ 3.3 Why is This So Common on Inductive Datasets?
In the last section, we covered when PPR can perform well on KGC and why. However, one remaining question is – why is this trend so pervasive on inductive datasets but rarely on their transductive counterparts? For example, on FB15k-237 [28], a transductive dataset, the Hits@10 of PPR is 2.7%. However, across eight different inductive versions of FB15k-237, the mean PPR performance is 32%. We find that this can be explained by the following two observations.
@@
But why does the mean SPD drastically change for negative samples but not positive? We find that this is due to the fact that entities in positive samples are more well-connected those those in negatives. In Table 1, we show the % difference in mean PPR score between positive and negative samples when controlling for the SPD on two inductive datasets.
Table 1: % Increase in in Pos over Neg PPR, controlling for SPD.
SPD Range WN18RR v4 FB15k-237 v4
@@
We can see that the PPR score is much greater for positive samples, even for higher values of SPD. Therefore, the SPD of positive samples are better able to “withstand" changes in the underlying graph better than negatives, as they typically contain additional shorter walks between samples. However, negatives samples are typically much less well-connected, so they are more affected by removing a portion of edges from the original graph.
(a) $\Delta\text{SPD}$ vs. PPR Hit@10
(b) PPR Hit@10 vs. Train Size
(c) PPR Hit@10 vs. Inference Size
@@ 4.1 Partition-Based Dataset Sampling
We note that the task of constructing inductive datasets from an existing graph $\mathcal{G}$ can be framed as a graph partitioning problem. Formally, we want to sample two non-overlapping partitions from the graph such that $\mathcal{G}_{\text{train}},\mathcal{G}_{\text{inf}}\in\mathcal{G}$. Given the analysis in Section 3.2, we hope to sample subgraphs such that $\Delta\text{SPD}(\mathcal{G}_{\text{train}})\approx\Delta\text{SPD}(\mathcal{G}_{\text{inf}})\approx\Delta\text{SPD}(\mathcal{G}$). But, how do we find subgraphs that satisfy this property? Intuitively, we want to sample each subgraph so that it’s removal has little effect on the initial graph’s structure. We give an example in Figure 5 where we sample two subgraphs from an existing graph. As we can see, even though the graph is partitioned, the relationship between entities in the same partition remain roughly the same before and after the partition. This is because there already exists little relationship between the two partitions in the original graph.
Table 2: $\Delta\text{SPD}$ and PPR Performance of New and Old Inductive Datasets vs. Transductive Parents.
Task Dataset PPR Hits@10 $\Delta\text{SPD}$
Trans. Ind. New Ind. Old Trans. Ind. New Ind. Old
(E) WN18RR 46.2 45.1 66.0 4.1 3.2 6.3
CoDEx-M 9.0 11.2 21.1 0.2 0.24 0.78
HetioNet 2.7 2.4 NA 0.55 0.29 NA
(E, R) FB15k-237 2.7 10.8 21.4 0.46 0.48 2.42
CoDEx-M 9.0 13.2 NA 0.20 0.42 NA

### v2

@@ 1 Introduction
Knowledge Graph Completion (KGC) attempts to predict unseen facts given an existing knowledge graph (KG). KGC has numerous applications including drug discovery [1], personalized medicine [2], and recommendations [3]. Traditionally, most research on KGC was focused on the transductive setting, where the same sets of entities and relations are seen during training and testing. Most methods [4, 5, 6] generally focus on learning embeddings for all entities and relations to facilitate the prediction of new facts.
In recent years, interest in KGC has shifted towards designing methods that can generalize to new entities or relations not seen during training. This task, known as “inductive KGC”, requires a method to train on a graph $\mathcal{G}_{\text{train}}$ and perform inference on a different graph $\mathcal{G}_{\text{inf}}$, where the inference graph contains either new entities, relations, or both. Because of this, methods for inductive KGC don’t rely on fixed embeddings for entities or relations, instead opting for more flexible techniques that can inductively learn representations based on a given graph [7, 8, 9]. To asses the ability of methods for this task, new datasets have been constructed that require methods to reason inductively. All inductive datasets [7, 10, 9] are constructed from existing transductive KGC datasets. This is done by sampling two graphs, one each for train and inference, which contain disjoint entities. Multiple methods [8, 11, 9] have reported tremendous promise on these newer benchmark datasets.
@@
This finding naturally motivates us to ask – why can PPR perform so well on existing inductive datasets? In Section 3 we discover that the high performance of PPR is due to how the inductive datasets are created from transductive datasets. Specifically, we observe that the current procedure creates graphs where the shortest path distance (SPD) between entities in positive test samples is much lower than the SPD between those in negative samples. This allows for a method like PPR, which gives a higher weight to shorter walks, to distinguish between the positive and negative samples based solely on the distance. To account for this problem, in Section 4 we propose a new strategy for sampling inductive datasets that uses graph partitioning to create the train and inference graphs. This allows us to sample subgraphs that retain the general properties of the original graph. We then demonstrate that this procedure can indeed create inductive datasets that greatly mitigates the ability of PPR. Lastly, we benchmark common inductive KGC methods on our newly constructed datasets. Our key contributions can be summarized as follows:
@@
  • Through empirical study, we find that the strong performance of PPR on inductive datasets is due to the current procedure for constructing inductive datasets, which allows for a shortcut to distinguish between positive and negative samples.
  • We propose a new strategy for sampling inductive KGC datasets from their transductive counterparts that uses graph partitioning. We show that our proposed method can substantially mitigate the shortcut. We then benchmark popular methods on our newly created datasets. Compared to the older datasets, methods tend to decrease in performance.
The remaining paper is structured as follows. In Section 2, we provide background on inductive KGC methods and datasets and PPR. In Section 3, we study in detail when and why PPR can perform well on inductive KGC. We then introduce our new strategy for creating inductive datasets in Section 4 and benchmark popular methods on these newly created datasets in Section 5.
@@ 2 Background and Related Work
Throughout this study we denote a knowledge graph as $\mathcal{G}=\{\mathcal{V},\mathcal{R},\mathcal{E}\}$ where $\mathcal{V}$ are the set of entities (i.e., nodes), $\mathcal{R}$ the set of relations (i.e., edge types), and $\mathcal{E}$ the set of edges (i.e., triples) of the form $(s,r,o)$ where $s$ and $o$ are entities and $r$ a relation. Lastly, we note that the task of KGC is formulated as the following where given a query $(s,r,*)$, we attempt to predict the correct entity *.
@@
Inductive KGC Datasets: In inductive KGC, we are given a training graph $\mathcal{G}_{\text{train}}=\{\mathcal{V}_{\text{train}},\mathcal{R}_{\text{train}},\mathcal{E}_{\text{train}}\}$ and an inference graph $\mathcal{G}_{\text{inf}}=\{\mathcal{V}_{\text{inf}},\mathcal{R}_{\text{inf}},\mathcal{E}_{\text{inf}}\}$. A method is trained on $\mathcal{G}_{\text{train}}$ and evaluated on $\mathcal{G}_{\text{inf}}$. Most datasets consider the setting that is only disjoint on the entities such that $\mathcal{R}_{\text{inf}}\subseteq\mathcal{R}_{\text{train}}$ and $\mathcal{E}_{\text{inf}}\cap\mathcal{E}_{\text{train}}=\emptyset$. This setting is referred to as the (E) setting. Another setting, which we denote as (E, R) further allows $\mathcal{R}_{\text{inf}}$ to contain relations not in $\mathcal{R}_{\text{train}}$.
@@
All existing inductive datasets are sampled from existing transductive datasets. Furthermore, the majority of inductive datasets [7, 9] are further created via the same procedure introduced by [7]. We now give a brief overview of this procedure. Given a transductive dataset, which we denote as $\mathcal{G}$, $k$ seed entities are randomly chosen from $\mathcal{G}$. The 2-hop neighborhood is then extracted around each individual seed entity. To prevent exponential growth, the number of neighbors sampled at any hop is capped at 50 for each seed entity. The resulting edges are then combined to create $\mathcal{G}_{\text{train}}$ and are subsequently removed from the original graph. The inference graph is then sampled in a similar manner from the resulting graph $\mathcal{G}\setminus\mathcal{G}_{\text{train}}$. One exception to this procedure are the ILPC datasets, introduced by [10], which instead sample $p\%$ of the nodes from $\mathcal{G}$ and use them to create $\mathcal{G}_{\text{train}}$. The rest of the nodes are then used to construct $\mathcal{G}_{\text{inf}}$. In practice, we find that both methods tend to produce similar graphs.
@@
Inductive KGC Methods: NeurlLP [13] and DRUM [14] consider combining the path representations of different length between both entities in a triple. However, since they explicitly consider each path, they are often limited to only considering paths of up to length 2 or 3. Conditional MPNNs [15] are a more efficient alternative to encoding higher-order path information. They work by conditioning the message passing mechanism on the known entity, allowing the implicit aggregation of all paths up to a length $L$ (which is equal to the number of GNN layers). The value of $L$ is typically $5/6$. Prominent examples include NBFNet [8] and RED-GNN [11]. More scalable alternatives have been proposed that prune the propagated messages [16, 17, 18]. NodePiece [19] is concerned with parameter efficiency, using an anchor-based approach to learn a more compact set of entity and relation representations. All previous methods assume that a fixed set of relations exist between the train and inference graphs. To account for new relations in inference, InGRAM [9] introduces the concept of a “relation graph”, which inductively encodes the representation of each relation. Note that we omit subgraph methods that have been to shown to prohibitively expensive [7, 20, 21, 22, 23] or those that require the use of external textual information [24, 25].
@@
where $W$ is the random walk matrix and $x_{s}$ is a one-hot vector at node $s$. Observe that the weight given to a walk decays with the increase in length due to $(1-\alpha)^{k}$. As such, the PPR score will often be higher for those nodes of shorter distance to $s$. For a KG, we obtain the PPR matrix by first converting the inference graph, $\mathcal{G}_{\text{inf}}$, to an undirected graph. This is common practice in KGC [27] whereby inverse relations are added to the graph. We further assume an edge weight of 1 for all edges. As such, the relations are completely disregarded when computing the PPR.
@@ 3.1 Personalized Pagerank (PPR) is a Strong Baseline for Inductive KGC
(a) % Increase in PPR performance from parent to inductive dataset
(b) $\Delta\text{SPD}$ vs. PPR Hits@10
Figure 2: (a) Percent increase in PPR performance on the inductive datasets as compared to their parent transductive dataset. For datasets with multiple splits (i.e., FB and WN) we take the mean performance across each. (b) Relationship between $\Delta\text{SPD}$ and performance. We observe a strong relationship among both transductive and inductive datasets.
We begin by obtaining the PPR matrix of the inference graph for use in evaluation. See Section 2 for more details on how this is done. Given this new graph, for a query $(s,r,*)$, we calculate the PPR score for all possible entities $o\in\mathcal{V}$. Using these scores, we can obtain the rank of the true entity for our query. We emphasize that PPR is (a) a non-learnable heuristic, (b) ignores the relations in the KG, and (c) has no basis in KG literature as a method to perform KGC.
In Figure 1, we show the performance when using the PPR versus the SOTA performance among supervised methods. The SOTA method is dataset dependent, with the specific methods listed in Appendix A.1. We split the datasets by transductive, (E), and (E, R) inductive. We include those datasets most often used in each task, comprising 25 in total. Please see Appendix B.1 for more on the datasets chosen. The full set of results can also be found in Appendix A.1.
We observe that on both types of inductive datasets, the PPR score does reasonably well, performing on average only 25-29% less than SOTA. This is surprising as PPR is both non-learnable and ignores the relational aspect of the KG. We also find that on some datasets like the WN or ILPC inductive splits, the performance nearly matches or exceeds the supervised performance. On the other hand, for the transductive datasets, the performance disparity is often much larger. Interestingly, we note that PPR still performs well on some transductive datasets, including the popular WN18RR. This tells us that this phenomenon is not necessarily unique to inductive datasets, but is most apparent there.
Furthermore, we find that for the inductive datasets, their PPR performance is much higher than their transductive parents. We detail this in Figure 2(a) where for different inductive datasets we see the percent increase in PPR performance from transductive to inductive. For example, on FB15k-237 the PPR Hits@10 is 2.7%, however the mean performance on the four FB (E) splits is 42.7%, representing a 1481% increase. We can see that the % increase on each inductive dataset is large, with the smallest being 42% on WN. This suggests that there is some change in the underlying inductive graphs that are causing the performance of PPR to increase. Lastly, we further explore other potential shortcuts in Appendix A.3, finding that PPR is by far the most severe.
@@ 3.2 When Does PPR Perform Well and Why?
We find that the answer lies in comparing the mean shortest path distance (SPD) between positive and negative samples. For a query $(s,r,*)$, let’s denote $o^{+}$ as the correct answer and $o^{-}\in\mathcal{V}^{-}(s,r)$ as the set of negative answers for the query. We compute the SPD on the inference graph for both the $o^{+}$ and $\mathcal{V}^{-}(s,r)$. This is repeated for each test sample $(s,q,*)\in\mathcal{E}_{\text{inf}}$. We then calculate the mean SPD across all positive and negative test samples, which we denote as $\text{SPD}^{+}$ and $\text{SPD}^{-}$, respectively. The different in mean SPD is correspondingly given by $\Delta\text{SPD}=\text{SPD}^{+}-\text{SPD}^{-}$. This tells us, on average, how much shorter the distance between entities in positive samples are relative to those in negatives.
In Figure 2(b) we show the relationship between $\Delta\text{SPD}$ and the Hits@10 when using PPR. The results are across 25 different datasets, for both inductive and transductive KGC. Calculating the Pearson correlation, we can observe it is 0.87, which indicates a strong relationship between the two metrics. Therefore, there exists a basic pattern to distinguish positive and negative samples in many KGC datasets. Put simply, the SPD between entities in positive test samples tend to be lower than that in negative samples. The larger the discrepancy between the mean distances, the better PPR can perform. This suggests that the PPR scores can exploit this pattern in the datasets, to achieve good performance, even while completely ignoring the relational information.
But, why can PPR exploit this pattern? As shown in Eq. (1), the PPR score between two nodes is the weighted sum of walks between them. Furthermore, walks of shorter length are weighed more heavily than those of longer length. Therefore, nodes with a shorter distance between them will be able to benefit from these highly-weighted walks, while those of larger distance will not. For example, when $\alpha=0.15$, the highest weight for a walk when $\text{SPD}=2$ is $0.72$ while when $\text{SPD}=5$ it is $0.44$. As such, the PPR score will invariably favor those node pairs with a lower SPD.
@@ 3.3 Why is This So Common on Inductive Datasets?
In the last section, we covered when PPR can perform well on KGC and why. However, one remaining question is – why is this trend so pervasive on inductive datasets but rarely on their transductive counterparts? For example, on FB15k-237, a transductive dataset, the Hits@10 of PPR is 2.7%. However, across eight different inductive versions of FB15k-237, the mean PPR performance is 32%. We find that this can be explained by the following two observations.
@@
Table 1: % Increase in Positive vs. Negative sample PPR, broken down by SPD.
SPD WN18RR v4 FB15k-237 v4
@@
But why does the mean SPD drastically change for negative samples but not positive? We find that entities in positive samples are more well-connected those those in negatives. In Table 1, we show the % difference in PPR score between positive and negative samples when controlling for the SPD on two inductive datasets.
We can see that the PPR score is much greater for positive samples, even for higher values of SPD. Therefore, the SPD of positive samples are better able to “withstand" changes in the underlying graph better than negatives, as they typically contain additional shorter walks between samples. However, negatives are typically much less well-connected, so they are more affected by removing a portion of edges from the original graph.
@@
(a) $\Delta\text{SPD}$ vs. PPR Hit@10
(b) Hits@10 vs. Train Size
(c) Hits@10 vs. Inference Size
@@ 4.1 Partition-Based Dataset Sampling
We note that the task of constructing inductive datasets from an existing graph $\mathcal{G}$ can be framed as a graph partitioning problem. Formally, we want to sample two non-overlapping partitions from the graph such that $\mathcal{G}_{\text{train}},\mathcal{G}_{\text{inf}}\subset\mathcal{G}$ and $\mathcal{G}_{\text{train}}\cap\mathcal{G}_{\text{inf}}=\emptyset$. Given the analysis in Section 3.2, we hope to sample subgraphs such that $\Delta\text{SPD}(\mathcal{G}_{\text{train}})\approx\Delta\text{SPD}(\mathcal{G}_{\text{inf}})\approx\Delta\text{SPD}(\mathcal{G}$). But, how do we find subgraphs that satisfy this property? Intuitively, we want to sample each subgraph so that it’s removal has little effect on the initial graph’s structure. We give an example in Figure 5 where we sample two subgraphs from an existing graph. As we can see, even though the graph is partitioned, the relationship between entities in the same partition remain roughly the same before and after the partition. This is because there already exists little relationship between the two partitions in the original graph.
@@
Table 2: Mean $\Delta\text{SPD}$ and PPR Hits@10 of new and old Inductive (“Ind.”) datasets vs. their Transductive (“Trans.”) parent. We highlight the value of the inductive dataset closer to the transductive as blue. For every case, the $\Delta\text{SPD}$ and PPR Hits@10 of the new inductive datasets are more aligned with the transductive dataset than the older inductive datasets.
Task Dataset Mean PPR Hits@10 Mean $\Delta\text{SPD}$
Trans. Ind. Old Ind. New Trans. Ind. Old Ind. New
(E) WN18RR 46.2 66.0 45.1 4.1 6.3 3.2
CoDEx-M 9.0 21.1 11.2 0.2 0.78 0.24
HetioNet 2.4 NA 2.7 0.55 NA 0.29
(E, R) FB15k-237 2.7 21.4 10.8 0.46 2.42 0.48
CoDEx-M 9.0 NA 13.2 0.20 NA 0.42